OpenStudy (anonymous):

Evaluate the following expression it using the values given find 2x-3y-x if x=-3,y=2, and z=-2

OpenStudy (xguardians):

= 2(-3) - 3(2) - (-3) = -6 - 6 + 3 = -12 + 3 = -9

OpenStudy (anonymous):

an RC circuit has an emf of 10 sin t volts, a resistance of 100 ohms, a capacitance of 0.005 farad, and no initial charge on the capacitor. Find (a) the charge on the capacitor at any time t and (b) the steady-state current.